# Caging in Time: A Framework for Robust Object Manipulation under Uncertainties and Limited Robot Perception

### v3

@@ 8.2.2 CBF and CLF
where $\mathcal{I}_{t}(i,j)$ is the probability at grid point $(i,j)$ in the discretized probability distribution, and $S(\mathbf{x})$ is the system entropy. The CLF is chosen as a weighted sum of the system’s energy and negative entropy to balance the goals of minimizing energy expenditure and maintaining a well-distributed probability state, ensuring both stability and robustness in the face of uncertainties.
@@
where $\frac{\partial h}{\partial\mathbf{x}}$ and $\frac{\partial V}{\partial\mathbf{x}}$ are the gradients of $h$ and $V$ with respect to the state $\mathbf{x}$, respectively.
In practice, these derivatives are approximated numerically due to the complexity of the expressions for $h$ and $V$.

### v4

@@ 9.2.2 CBF and CLF
where $\mathcal{I}_{t}(i,j)$ is the probability at the grid point $(i,j)$ in the discretized probability distribution, and $S(\mathbf{x})$ is the system entropy. The CLF is chosen as a weighted sum of the system’s energy and negative entropy to balance the goals of minimizing energy expenditure and maintaining a well-distributed probability state, ensuring both stability and robustness in the face of uncertainties.
@@
where $\frac{\partial h}{\partial\mathbf{x}}$ and $\frac{\partial V}{\partial\mathbf{x}}$ are the gradients of $h$ and $V$ with respect to the state $\mathbf{x}$, respectively. In practice, these derivatives are approximated numerically due to the complexity of the expressions for $h$ and $V$.